OpenStudy (anonymous):

lil123 would 7p^3r - 5t^3s - 31p^2r - 10s^3t^3 = 7p^2r(p - 3) 5t^2s(t-2s^2t)?

OpenStudy (lilg132):

7p ^2r * p = 7p^3r 7p^2r * -3 = -21p^2r That is wrong

OpenStudy (lilg132):

31 is a prime number so i would leave the numbers inside the bracket instead of outside

OpenStudy (lilg132):

so have for the first bracket

OpenStudy (lilg132):

p^2r(7p - 31)

OpenStudy (anonymous):

i apologize the 31 is supposed to be 21

OpenStudy (lilg132):

5t^2s(t-2s^2t) that would be 5t^2s * t = 5t^3s which is correct 5t^2s * -2s^2t = 10s^3t^3 which is correct also my bad your right

OpenStudy (lilg132):

second equation should be = -10s^3 t^3 my mistake